# A General-Purpose Multimodal Foundation Model for Dermatology

### v1

@@ Methods
Most datasets used in this study are publicly available. These datasets used for skin lesion diagnosis and segmentation tasks can be accessed through various repositories. The ISIC archive141414isic-archive.com hosts several datasets, including MSKCC and HIBA. Other widely used benchmark datasets are available through their respective portals: BCN20000151515figshare.com/articles/journal_contribution/BCN20000_Dermoscopic_Lesions_in_the_Wild/24140028/1, PAD-UFES-20161616kaggle.com/datasets/mahdavi1202/skin-cancer, DDI171717ddi-dataset.github.io/index.html, Derm7pt181818derm.cs.sfu.ca/Welcome.html, ISIC2024191919kaggle.com/competitions/isic-2024-challenge, Med-Node202020kaggle.com/datasets/prabhavsanga/med-node, DermNet212121kaggle.com/datasets/shubhamgoel27/dermnet, WSI222222portal.gdc.cancer.gov/projects/TCGA-SKCM, PATCH16232323heidata.uni-heidelberg.de/dataset.xhtml?persistentId=doi:10.11588/data/7QCR8S, ISIC2018_task1 and HAM1000242424challenge.isic-archive.com/data/, SDDI1252525api.isic-archive.com/collections/328/, and PH2262626fc.up.pt/addi/ph2%20database.html. Access to in-house datasets is restricted due to patient privacy considerations. These include MMT for dermoscopic and clinical image pretraining and downstream multi-skin condition classification, NSSI for sequential dermoscopic image pretraining, ACEMID_path for dermatopathology pretraining, Edu1 and Edu2 for clinical image pretraining, SDDI2 for lesion change detection, SDDI_Alfred for reader study 1 (early-melanoma detection), and the TBP data from MYM and HOP studies for all TBP-based pretraining and evaluation. Researchers interested in accessing these datasets should direct their requests to the corresponding author. Requests will be evaluated according to institutional and departmental policies to ensure compliance with intellectual property rights and patient privacy obligations. The availability of these data may be subject to additional restrictions or requirements.
@@
S.Y., Z.Y., H.K., H.P.S, and Z.G. conceived the study and designed experiments. S.Y., C.V.A., M.H., L.Y., H.K., V.M., M.J., H.P.S., and Z.G. contributed to data acquisition, preprocessing, and organization. S.Y. performed model development and pretraining. S.Y., Z.Y., Z.W., L.Y., H.K., and P.T. contributed to downstream task evaluation. S.Y., Z.Y., C.V.A., and V.M. performed experimental analyses regarding metastasis prediction and prognosis. S.Y. and Z.Y. performed experimental analyses regarding TBP-based screening. Z.W. and S.Y. performed experimental analysis regarding lesion segmentation. S.Y. and L.Y. performed experimental analysis regarding dermatopathology image analysis. S.Y., H.K., and P.T. contributed to the human-AI collaboration reader study experimental analysis. S.Y. also contributed to all remaining tasks of experimental analysis. H.K. and P.T. developed the web-based reader platforms and conducted the human-AI reader study. V.M. and Z.Y. contributed to the early melanoma detection reader study data. H.K., P.T., C.V.A., C.P., V.M., M.J., and H.P.S. contributed clinical inputs to this research. G.T., V.T., A.B.N., D.P., P.B., and S.S. provided computing resources and data management. All authors contributed to the drafting and revising of the manuscript.
@@
Model Accuracy (ACC)
PanDerm 0.932
BioMedGPT 0.866
MedViT 0.723
BiomedCLIP 0.719
Extended Data Table 30: Model performance comparison with GMAI on skin cancer classification using HAM10000 dataset. Accuracy (ACC) is reported for different models, including the generalist medical model BioMedGPT. PanDerm achieves the highest accuracy. Performance of baseline models are as reported in [60].
@@
Extended Data Table 32: PanDerm hyperparameters used in finetuning. Single 49GB NVIDIA 6000Ada GPU was used for downstream finetuning.

### v2

@@ Methods
Most datasets used in this study are publicly available. These datasets used for skin lesion diagnosis and segmentation tasks can be accessed through various repositories. The ISIC archive141414isic-archive.com hosts several datasets, including MSKCC and HIBA. Other widely used benchmark datasets are available through their respective portals: BCN20000151515figshare.com/articles/journal_contribution/BCN20000_Dermoscopic_Lesions_in_the_Wild/24140028/1, PAD-UFES-20161616kaggle.com/datasets/mahdavi1202/skin-cancer, DDI171717ddi-dataset.github.io/index.html, Derm7pt181818derm.cs.sfu.ca/Welcome.html, ISIC2024191919kaggle.com/competitions/isic-2024-challenge, Med-Node202020kaggle.com/datasets/prabhavsanga/med-node, DermNet212121kaggle.com/datasets/shubhamgoel27/dermnet, WSI222222portal.gdc.cancer.gov/projects/TCGA-SKCM, PATCH16232323heidata.uni-heidelberg.de/dataset.xhtml?persistentId=doi:10.11588/data/7QCR8S, ISIC2018_task1 and HAM1000242424challenge.isic-archive.com/data/, SDDI1252525api.isic-archive.com/collections/328/, PH2262626fc.up.pt/addi/ph2%20database.html, and SD-128 272727huggingface.co/datasets/resyhgerwshshgdfghsdfgh/SD-198. Access to in-house datasets is restricted due to patient privacy considerations. These include MMT for dermoscopic and clinical image pretraining and downstream multi-skin condition classification, NSSI for sequential dermoscopic image pretraining, ACEMID_path for dermatopathology pretraining, Edu1 and Edu2 for clinical image pretraining, SDDI2 for lesion change detection, SDDI_Alfred for reader study 1 (early-melanoma detection), and the TBP data from MYM and HOP studies for all TBP-based pretraining and evaluation. Researchers interested in accessing these datasets should direct their requests to the corresponding author. Requests will be evaluated according to institutional and departmental policies to ensure compliance with intellectual property rights and patient privacy obligations. The availability of these data may be subject to additional restrictions or requirements.
@@
S.Y., H.K., and Z.G. conceived the study and designed experiments. S.Y., C.V.A., M.H., L.Y., L.J., H.K., V.M., M.J., H.P.S., and Z.G. contributed to data acquisition, preprocessing, and organization. S.Y. performed model development and pretraining. S.Y., Z.Y., Z.W., L.Y., H.K., and P.T. contributed to downstream task evaluation. S.Y., Z.Y., C.V.A., E.M., P.G, J.B., and V.M. contributed to the metastasis prediction and prognosis tasks. S.Y. and Z.Y. performed experimental analyses regarding TBP-based screening. Z.W. and S.Y. performed experimental analysis regarding lesion segmentation. S.Y., L.Y., and P.F. contributed to dermatopathology image analysis tasks. S.Y., H.K., P.T., V.M., and H.P.S. contributed to the two human-AI collaboration reader studies. S.Y. also contributed to all remaining tasks of experimental analysis. H.K. and P.T. developed the web-based reader platforms and conducted the human-AI reader study. V.M., J.N., and Z.Y. contributed to the early melanoma detection reader study data. H.K., P.T., C.V.A., C.P., V.M., M.J., and H.P.S. contributed clinical inputs to this research. G.T., V.T., A.B.N., D.P., P.B., and S.S. provided computing resources and data management. All authors contributed to the drafting and revising of the manuscript.
Acknowledgements
We thank the following readers who participated in the reader study 3: Ali Abid, Ana Brkic, Marija Buljan, Giuliana Carlos Minano, Charmaine Chamberlin, Mary ElSharouni, Jon Fulgencio, Klara Gacina, Emilia Garcia, Victor Giang, Prue Gramp, Claudia Huang, Emma Karlsen, Damian Kostner, Jack Lai, Francesco Leo, Sophie Lim, Elisabetta Magnaterra, Isabel Matheus, Rebeca Mielgo Salvador, Christoph Mueller, Nicholas Muller, Jennifer Nguyen, Maria Onteniente, Nisal Punchihewa, Polina Reztsova, Belen Rosales Trujillo, Cliff Rosendahl, Helmut Schaider, Christine Shen, Juliet Smith, Corey Stone, Tom Sun, Deborah Toledo Flores, Cathleen Ventura, Anna Wolber, and Mabel Yan.
@@ Acknowledgements
Extended Data Figure 3: Early melanoma detection results (reader study 1): Comparing PanDerm to 12 clinicians (7 experienced dermatologists, 5 dermatology residents). X-axis: 89 melanoma lesion IDs; Y-axis: lesion image sequence length. Points on the histogram represent the initial time points of correct melanoma diagnoses. Points below y=0 correspond to melanoma lesions undetected throughout the sequence
Extended Data Figure 4: Performance of PanDerm in Human-AI Collaborative 7 Skin Cancer Diagnosis using Dermoscopic Images. a. Overview of the reader study: 41 users participated, answering a total of 3,200 questions. The maximum, median, and minimum number of questions answered per user were 1,040, 40, and 20, respectively. b. Comparison of diagnostic accuracy without support and with PanDerm support (P $<$ 0.001; two-sided paired t-test; n = 41 readers). c. Comparison of diagnostic accuracy without support and with PanDerm support, grouped by experience level: Low (n = 11 readers), Medium (n = 21 readers), and High (n = 9 readers). d. Comparison of diagnostic accuracy without support and with PanDerm support, grouped by diagnostic classes. Abbreviations: MEL, melanoma; BCC, basal cell carcinoma; AKIEC, actinic keratosis/intraepidermal carcinoma; BKL, benign keratinocytic lesion; NV, melanocytic nevus; DF, dermatofibroma; VASC, vascular lesion.
Extended Data Figure 5: Sunburst plot of standard ontology on SD-128 dataset.
Extended Data Figure 6: Participant specialties distribution for reader study 3.
Extended Data Figure 7: Participant career distribution for reader study 3.
Extended Data Figure 8: Participant experience levels by years for reader study 3.
@@
Extended Data Figure 12: Longitudinal dermoscopic image-based lesion change detection using PanDerm. For comparing subtle changes in paired lesions during short-term follow-up (e.g., 3 months), images undergo dark corner detection and removal, skin inpainting, registration, and lesion segmentation. This allows models to focus on subtle differences between lesions at different time points.
Characteristic n (%) or Mean ± SD
Total Participants
Total 193
Age
Age (years) 52.2 ± 12.0
Sex
Female 82 (42.5%)
Male 111 (58.5%)
Ancestry/Ethnicity
European/British 164 (85.0%)
Mixed/Other 29 (15.0%)
Skin Phototype
Burns easily, tans slightly 52 (26.9%)
Burns moderately, tans gradually 123 (63.7%)
Rarely burns, tans well 18 (9.3%)
Innate Skin Color
Fair/Type I 144 (74.6%)
Medium/Type II 48 (24.9%)
Olive/Type III 1 (0.5%)
Extended Data Table 1: Demographic characteristics of MYM cohort.
Characteristic N (%) or Mean ± SD
Sample Size
Total 314
Age
Age (years) 56.1 ± 12.8
Sex
Male 120 (38.2%)
Female 194 (61.8%)
Ancestry
European/British 297 (94.6%)
Other/Mixed 17 (5.4%)
Birth Place
Australia 268 (85.4%)
Overseas 46 (14.6%)
Melanoma History
Yes 304 (96.8%)
No 10 (3.2%)
Non-melanoma Skin Cancer
Yes 195 (62.1%)
No 119 (37.9%)
Innate Skin Color
Fair/Type I 269 (85.7%)
Medium/Type II 44 (14.0%)
Olive/Type III 1 (0.3%)
Extended Data Table 2: Demographic and clinical characteristics of HOP cohort.
Characteristic N (%)
Age
$>$ 60 years 190 (51.4)
$\leq$ 60 years 180 (48.6)
Sex
Female 172 (47.3)
Male 198 (53.5)
Stage at Diagnosis
Stage I 261 (70.5)
Stage II 45 (12.2)
Stage III 61 (16.5)
Stage IV 3 (0.8)
T Classification
T1a 193 (59.2)
T1b 28 (8.6)
T2a 56 (18.6)
T2b 13 (4.0)
T3a 21 (6.4)
T3b 11 (3.4)
T4a 7 (2.1)
T4b 42 (13.2)
Unknown 1 (0.3)
SLNB Status
Not Performed 265 (71.6)
Positive 40 (10.8)
Negative 65 (17.6)
N Classification
N1 40 (10.8)
N2 14 (3.8)
N3 7 (1.8)
Unknown 2 (0.5)
Extended Data Table 3: Demographic characteristics of the CombinMel dataset.
Characteristic Melanoma N (%) Benign N (%)
Gender
Male 48 (53.9) 36 (40.0)
Female 41 (46.1) 54 (60.0)
Previous melanoma 67 (75.2) 57 (63.3)
Previous NMSC 44 (49.4) 20 (22.2)
Family history of melanoma 49 (55.0) 38 (42.2)
Naevi count
<20 0 (0.0) 0 (0.0)
20-50 3 (3.4) 1 (1.1)
50-100 27 (30.3) 15 (16.7)
100-200 14 (15.7) 33 (36.7)
200-500 42 (47.2) 39 (43.3)
>500 3 (3.4) 2 (2.2)
Atypical Naevi count
<5 19 (21.3) 22 (24.4)
5-15 28 (31.5) 29 (32.2)
>15 42 (47.2) 39 (43.0)
Fitzpatrick Skin Phototype
I 12 (13.4) 11 (12.2)
II 42 (47.1) 47 (52.2)
III 28 (31.4) 32 (35.6)
IV 0 (0.0) 0 (0.0)
V 7 (7.8) 0 (0.0)
Location
Head and neck 7 (7.8) 7 (7.8)
Chest 9 (10.1) 9 (10.0)
Abdomen 7 (7.8) 14 (15.6)
Back 21 (23.5) 23 (25.6)
Upper limb 19 (21.3) 13 (14.4)
Lower limb 26 (29.2) 24 (26.7)
Histopathology (Melanoma)
In situ unspecified 16 (18.0)
In situ SSM 28 (31.4)
Lentigo maligna 11 (12.3)
Invasive SSM 32 (36.0)
Invasive LMM 2 (2.2)
Histopathology (Benign)
Dysplastic naevus 36 (40.0)
Compound naevus 25 (27.8)
Junctional naevus 17 (18.9)
Intradermal naevus 8 (8.9)
Other 4 (4.4)
Extended Data Table 4: Demographic characteristics of the SDDI_Alfred dataset.
Characteristic N (%)
Total Participants 1254
Age (years)
Mean 47.22
Median 46
Range 11-88
Sex
Male 602 (48.0)
Female 652 (52.0)
Melanoma History
Personal history of melanoma 589 (47.0)
No personal history of melanoma 665 (53.0)
Facultative Skin Colour (n=1245)
Fair/Type 1 635 (51.0)
Medium/Type 2 510 (41.0)
Olive/Type 3 100 (8.0)
Self-reported Ancestry*
British/Irish 963
West/Northern European 228
No data or other (not specified) 104
Southern European 36
Eastern European 23
Asian 18
Middle Eastern 8
Other (specified) 8
Indigenous Australian 4
Pacific Islander 4
Extended Data Table 5: Demographic characteristics of the NSSI dataset.
* Multiple ancestries could be reported by each participant
Characteristic N (%)
Total Participants 54
Study Sites
QLD 1 - PAH 26 (48.1)
NSW 3 - MIA 28 (51.9)
Age (years)
Range 19-75
Mean 53.4
Gender
Male 31 (57.4)
Female 23 (42.6)
Melanoma Risk Group
Very High 44 (81.5)
High 8 (14.8)
Low/Average 1 (1.9)
Unknown 1 (1.9)
Primary Diagnosis (Pathologist 1)
Total Naevi 37 (68.5)
   Common/Dermal/Congenital 14 (25.9)
   Dysplastic Compound 22 (40.7)
   Dysplastic Junctional 8 (14.8)
Total Melanomas 13 (24.1)
   In Situ 11 (20.4)
   Thin Invasive 2 (3.7)
Other Lesions 4 (7.4)
Number of Lesions per Patient
1 36 (66.7)
2 12 (22.2)
3 2 (3.7)
4 2 (3.7)
5 2 (3.7)
Extended Data Table 6: Demographic characteristics of the ACEMID_pilot_study dataset.
Major Category Classes
Skin Cancer and High-Risk Lesions (13) Basal Cell Carcinoma - Nodular, Basal Cell Carcinoma - Superficial, Basal Cell Carcinoma - Pigmented, Basal Cell Carcinoma - Recurrent, Melanoma - Nodular, Lentigo Maligna, Squamous Cell Carcinoma - Common, Squamous Cell Carcinoma - Bowen’s, Keratoacanthoma, Actinic Keratosis - Common, Actinic Keratosis - Hypertrophic, Actinic Keratosis - Pigmented, Actinic Cheilitis
Melanocytic Lesions (22) Compound Nevus, Junctional Nevus, Dermal Nevus, Acral Parallel Pattern Nevus, Acral Lattice Pattern Nevus, Acral Untyped Pattern Nevus, Blue Nevus, Spitz Nevus, Reed Nevus, Congenital Nevus, Halo Nevus, Agminated Nevus, Atypical Nevus, Solar Lentigo, Lentigo Simplex, Ink-spot Lentigo, Ephelides, Melanosis, En Cocarde Nevus, Lentiginous Nevus, Subungual Melanocytic Nevus, Nevus with Regression
Common Benign Growths (13) Seborrheic Keratosis - Common, Seborrheic Keratosis - Inflamed, Seborrheic Keratosis - Pigmented, Dermatofibroma, Epidermal Cyst, Sebaceous Hyperplasia, Skin Tag, Fibrous Papule of Face, Cutaneous Horn, Comedone, Accessory Nipple, Myxoid Cyst, Chondrodermatitis
Inflammatory Conditions (8) Psoriasis, Eczema, Dermatitis, Folliculitis, Granuloma Annulare, Lichenoid Dermatitis, Porokeratosis, Hypertrophic Lichen Planus
Vascular and Infectious Lesions (10) Angiokeratoma, Angioma, Telangiectasia, Hematoma - Common, Hematoma - Subcorneal, Hematoma - Subungual, Molluscum Contagiosum, Wart, Vascular Malformation, Pyogenic Granuloma
Others (8) Post-inflammatory Changes, Nail Dystrophy, Hypertrophic Scar, Atrophic Scar, Excoriation, Traumatic Changes, Striae, Chemical Burn
Extended Data Table 7: The taxonomy of 74 dermatological conditions covered in MMT-74 dataset.
@@
Study Mean Recall Recall_Mel Recall_BCC
Tschandl et al. 2019 [75] 0.777 0.614 0.796
PanDerm 0.804 0.877 0.860
Extended Data Table 30: Performance comparison with prior Human-AI collaboration work in HAM10000. Comparison of model performance between PanDerm and [75] on HAM10000 dataset. PanDerm shows improvements in overall recall (+2.72%) and particularly significant gains for critical cancers like melanoma (+26.32%) and BCC (+6.42%).
@@
Modality Tasks Dataset Metric ResNet50 BioMedGPT PanDerm
TBP Melanoma detection ISIC2024 AUROC 0.8023∗∗ 0.6840∗∗∗ 0.893
(2 classes) (0.7327-0.8723) (0.5670-0.8009) (0.839-0.940)
Dermoscopic Metastasis Prediction CombinMel AUROC 0.9378 0.7774∗∗∗ 0.964
images (2 classes) (0.8979-0.9618) (0.7094-0.8453) (0.937-0.991)
Dermoscopic Skin cancer classification HAM_clean Weighted 0.8606∗∗∗ 0.8309∗∗∗ 0.926
images (7 classes) F1 (0.8411-0.8771) (0.8142-0.8476) (0.912-0.940)
Clinical General(common) skin DermNet Weighted 0.4181∗∗∗ 0.3086∗∗∗ 0.619
images condition classification F1 (0.4039-0.4323) (0.2952-0.3220) (0.603-0.634)
(23 classes)
Clinical General skin condition Fitzpatrick1 Weighted 0.2578∗∗∗ 0.1350∗∗∗ 0.4817
images classification 7K_clean F1 (0.2410-0.2747) (0.1225-0.1475) (0.4628-0.5005)
(115 classes)
Dermato- Tumor classification PATCH16 Weighted 0.8304∗∗∗ 0.7066∗∗∗ 0.903
pathology (16 classes) F1 (0.8242-0.8365) (0.6991-0.7142) (0.898-0.908)
Extended Data Table 38: Performance comparison across different dermatological tasks and modalities. Models are evaluated on various dermatological tasks spanning TBP, dermoscopic images, clinical photographs, and dermatopathology. Evaluation metrics include Area Under the Receiver Operating Characteristic curve (AUROC) for binary classification tasks and Weighted F1 score for multi-class classification. Performance is reported with 95% confidence intervals in parentheses. The best performance for each task is bolded. ${}^{***}p<0.001$ compared to PanDerm.
Model Accuracy Weighted F1
BioMedGPT (finetune) 0.7394 0.7740
BioMedGPT (Linear probe) 0.7149 0.6816
PanDerm (Finetune) 0.8538 0.8538
PanDerm (Linear probe) 0.8306 0.8252
Extended Data Table 39: Performance comparison between BioMedGPT and PanDerm models on HAM10000. Models are evaluated using both fine-tuning and linear probing (LP) approaches. Performance is reported using Accuracy and Weighted F1 score metrics. Best performance for each metric is bolded.
@@
Extended Data Table 41: PanDerm hyperparameters used in finetuning. A single 49GB NVIDIA 6000Ada GPU was used for downstream finetuning.